Logiweb codex of factorial

Up Help factorial, ∗!,

factorial

[factorial →^{pyk} “factorial”]

∗!

[x! →^{val} If(x ≈ 0, 1, x · x − 1!)]

[x! →^{tex} “#1.!”]

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-26.UTC:10:24:54.494911 = MJD-53516.TAI:10:25:26.494911 = LGT-4623819926494911e-6